# MegaSynth: Scaling Up 3D Scene Reconstruction with Synthesized Data

### v1

@@ Abstract
We propose scaling up 3D scene reconstruction by training with synthesized data. At the core of our work is MegaSynth, a procedurally generated 3D dataset comprising 700K scenes—over 50 times larger than the prior real dataset DL3DV—dramatically scaling the training data. To enable scalable data generation, our key idea is eliminating semantic information, removing the need to model complex semantic priors such as object affordances and scene composition. Instead, we model scenes with basic spatial structures and geometry primitives, offering scalability. Besides, we control data complexity to facilitate training while loosely aligning it with real-world data distribution to benefit real-world generalization. We explore training LRMs with both MegaSynth and available real data. Experiment results show that joint training or pre-training with MegaSynth improves reconstruction quality by 1.2 to 1.8 dB PSNR across diverse image domains. Moreover, models trained solely on MegaSynth perform comparably to those trained on real data, underscoring the low-level nature of 3D reconstruction. Additionally, we provide an in-depth analysis of MegaSynth’s properties for enhancing model capability, training stability, and generalization.
@@ 1 Introduction
In this work, we propose scaling up training data for scene-level reconstruction by using synthesized data. Our key idea is to eliminate the reliance on semantic information in data generation, by constructing scenes with non-semantic shape primitives arranged within basic spatial structures. This approach is motivated by our insight that scene semantics play a minimal role in multi-view reconstruction, as evidenced by the success of traditional non-semantic methods such as COLMAP [57], MVS [59], NeRF [48], and the emerging non-semantic properties of recent feed-forward models [29, 9, 76]. Unlike prior scene generation methods, which aim to replicate real-world scene distributions [54, 63, 16, 72, 53] and are thus constrained by the complexity of modeling semantics, such as object affordances, our approach bypasses these challenges. This simplification enables highly scalable and efficient data generation.
@@
We generate the MegaSynth dataset, comprising 700K scenes. MegaSynth is over 50 times larger than the real dataset DL3DV and significantly scales up training data for LRMs. We utilize MegaSynth to train feed-forward LRMs [85, 89] jointly with DL3DV. Our experiments show a 1.2 to 1.8 dB PSNR gain across diverse test datasets and image resolutions, highlighting the synergy between synthesized and real data. Synthesized data excels in scale and provides rich metadata, such as geometry supervision, enabling models to develop a general geometric understanding beyond rendering supervision. Meanwhile, small-scale real data further sharpens the model. Notably, models trained solely on MegaSynth, regardless of the number of input views, perform comparably to those trained on real data. This result supports our hypothesis that 3D reconstruction is predominantly a low-level vision task, with minimal reliance on semantics.
@@ 4 Synthesizing the MegaSynth Dataset
4.1 Overview
We synthesize MegaSynth using a procedural generation method, as illustrated in Fig. 2. The process involves: i) generating a scene floor plan, including the scene size and object instance box locations, ii) instantiating object geometries with random textures, and iii) randomizing the lighting. During this process, we eliminate high-level scene semantics, e.g., category-related object affordance. We only keep the low-level structural and geometric features of scenes.
@@ 4.2 Scene Floor Plan
Without loss of generality, we plan the scene as a cube box and populate it with objects represented by 3D bounding boxes. We randomize the 3D aspect ratio and size of scenes. We design multiple object box categories to simulate real-world scene geometry structures (visualized as boxes in Fig. 2 with different colors). For example, large object boxes tend to be placed near the ground while small object boxes have more flexible placement options. We parameterize the size, location, and number of each object type, specifying each parameter as a range. This allows us to introduce randomness to improve diversity. Further details of the object box categories and their attribute sampling ranges are provided in the Appendix A.
@@ 4.4 Lighting
Luminous objects and light bulbs. We randomly turn some objects and axis-aligned sticks in the scene as lights to simulate real-world lighting. We also randomly place light bulbs in the scene to increase diversity. The intensity of object light can be sampled as large values to simulate lighting in dark environments.
@@ 5 Learning 3D Reconstruction on MegaSynth
In this section, we discuss how we utilize our synthesized MegaSynth (Sec. 4) to train a feed-forward reconstruction model (i.e., the LRM-based model illustrated in Sec. 3). To reach the goal, we first construct the training data by carefully sampling the camera distribution and rendering the images (Sec. 5.1). We then train our model with a mixed-data training strategy (Sec. 5.2) with rendering loss and geometry loss (Sec. 5.3). The details of the training process can be found in the Appendix B.
@@ 6 Experiments
Model Training Dataset PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
@@
3DGS [35] N.A. (Per-scene Optimization) 5.2 min 24.27 0.817 0.166 20.67 0.672 0.293 16.46 0.458 0.405
Long-LRM [89] DL3DV 0.12 sec 24.18 0.812 0.173 23.41 0.790 0.210 19.68 0.569 0.312
Long-LRM (ours) DL3DV + MegaSynth 25.44 0.853 0.136 25.01 0.836 0.164 20.86 0.652 0.249
GS-LRM [85] DL3DV 0.11 sec 24.60 0.824 0.161 23.89 0.806 0.195 19.93 0.601 0.289
GS-LRM (ours) DL3DV + MegaSynth 25.75 0.859 0.130 25.46 0.846 0.154 21.19 0.672 0.235
@@
3DGS [35] N.A. (Per-scene Optimization) 6.4 min 23.26 0.778 0.206 21.75 0.690 0.294 16.06 0.436 0.421
Long-LRM [89] DL3DV 0.35 sec 23.71 0.779 0.236 22.51 0.767 0.267 18.61 0.465 0.421
Long-LRM (ours) DL3DV + MegaSynth 25.14 0.828 0.186 24.26 0.817 0.210 19.84 0.555 0.339
Table 1: Evaluation results against baseline methods. We report results resolution of 128 and 256. For resolution 256, we only report results of Long-LRM as transformer-based GS-LRM is too slow. Our models are pre-trained on MegaSynth and then tuned on DL3DV.
In this section, we describe the experimental setting and present evaluation results. Due to the space limit, implementation and training details are in the Appendix B.
@@ 6.1 Datasets
Hypersim [54] is a synthetic 3D indoor scene dataset with ultra photo-realistic renderings, aimed at testing the generalization capability to out-of-distribution indoor scenes. Hypersim is challenging due to its complicated geometry, extreme lighting conditions, and large camera baselines. We use a test set composed of 302 scenes.
@@ 6.2 Evaluation and Baselines
We use $32$ views as input and use $32$ target novel view images for evaluation. The input and target views are non-overlap and are evenly sampled. We compare with three baseline methods:
@@
Additionally, we note that comparing with more advanced 3DGS methods is not the focus our work, as our target is scaling up training data for improving feed-forward methods. 3DGS is only used to verify our overall performance, and our contributions can be directly ablated by comparing with LRMs trained without our data.
@@ 6.3 Results
Table 1 presents our results, demonstrating that training with both DL3DV and our MegaSynth dataset significantly improves model performance, with PSNR gains ranging from 1.2 to 1.8 dB. This improvement is consistent across model architectures (GS-LRM and Long-LRM), testing data (both in-domain DL3DV and out-of-domain datasets), image resolutions, and evaluation metrics, highlighting the effectiveness of our synthesized MegaSynth in enhancing the reconstruction quality of LRMs for wide-coverage scenes. Fig. 3 and Fig. 4 visually compare the reconstruction results for models with and without MegaSynth. We observe remarkable improvements in scenes with complicated scene structures, geoemtry, material and lighting, aligning with data generation designs (Sec. 4). Our approach also achieves substantially better results than the optimization-based 3DGS method while offering much faster inference speeds (e.g., over 2000 times speed-up from 5 minutes to 0.1 seconds).
@@ 6.4 Ablation Studies
In this section, we examine the impact of synthesized data quality, quantity, and properties on real-data performance, along with training paradigms for utilizing synthesized data. Without additional specification, the default experimental setup in analysis is the resolution-128 GS-LRM with pre-training + fine-tuning training protocol.
@@
DL3DV-Test-Indoor DL3DV-Test-Outdoor
Data PSNR SSIM LPIPS PSNR SSIM LPIPS
DL3DV 25.41 0.853 0.150 23.09 0.771 0.183
DL3DV + MegaSynth 26.75 0.890 0.116 23.89 0.803 0.157
Table 3: Performance gains on indoor and outdoor test data. Results are from 128-resolution GS-LRM. The indoor test is our held-out set and the outdoor test is the DL3DV benchmark. Data details in Sec. 6.1.
Hypersim DL3DV-indoor
PSNR SSIM LPIPS PSNR SSIM LPIPS
Real-only 23.89 0.806 0.195 25.41 0.853 0.149
MegaSynth-only 21.50 0.719 0.272 23.02 0.782 0.217
Joint Training 25.33 0.844 0.157 26.54 0.885 0.121
Pre-training + Fine-tuning 25.46 0.846 0.154 26.75 0.890 0.116
Table 4: Ablation study on the training framework to leverage synthesized data. Results are reported on Hypersim and DL3DV indoor test subset with GS-LRM.
Table 5: Visual comparison between models trained with only MegaSynth and only real data. We include two failure cases of the MegaSynth-only model and highlight the failure regions. Results are reported with resolution 128 on Hypersim.
MegaSynth
@@
w. Geomtry 23.24 0.684 0.280
w. Material 23.07 0.663 0.336
w. Lighting 22.65 0.660 0.305
Table 6: Analysis of the synthesized data control. Results are reported with resolution 128. The model is only trained on DL3DV to measure the distribution distance.
Table 7: Analysis of the NVS quality with different input view numbers. Results are reported with resolution 128 on Hypersim.
@@ 6.5 Analysis
We analyze the data control strategy and MegaSynth’s effectiveness with varying input views.
Analysis of data control. We evaluate the quality of MegaSynth and its alignment with real data by applying the GS-LRM model, trained on real-world DL3DV data, to our generated MegaSynth. As shown in Table 7, the reconstruction performance is reasonable with a 23 dB PSNR, indicating good alignment with real data. Notably, by controlling geometry, material, and lighting (introduced in Sec. 4) step by step, the performance slightly decreases, suggesting an increased data complexity.

### v2

@@ Abstract
We propose scaling up 3D scene reconstruction by training with synthesized data. At the core of our work is MegaSynth, a procedurally generated 3D dataset comprising 700K scenes—over 50 times larger than the prior real dataset DL3DV—dramatically scaling the training data. To enable scalable data generation, our key idea is eliminating semantic information, removing the need to model complex semantic priors such as object affordances and scene composition. Instead, we model scenes with basic spatial structures and geometry primitives, offering scalability. Besides, we control data complexity to facilitate training while loosely aligning it with real-world data distribution to benefit real-world generalization. We explore training LRMs with both MegaSynth and available real data. Experiment results show that joint training or pre-training with MegaSynth improves reconstruction quality by 1.2 to 1.8 dB PSNR across diverse image domains. Moreover, models trained solely on MegaSynth perform comparably to those trained on real data, underscoring the low-level nature of 3D reconstruction. Additionally, we provide an in-depth analysis of MegaSynth’s properties for enhancing model capability, training stability, and generalization, as well as application to other tasks.
@@ 1 Introduction
In this work, we propose scaling up training data for scene-level reconstruction by using synthesized data. Our key idea is to eliminate the reliance on semantic information in data generation, by constructing scenes with non-semantic shape primitives arranged within basic spatial structures. This approach is motivated by our insight that scene semantics play a minimal role in multi-view reconstruction, as evidenced by the success of traditional non-semantic methods such as COLMAP [60], MVS [62], NeRF [50], and the emerging non-semantic properties of recent feed-forward models [31, 9, 79]. Unlike prior scene generation methods, which aim to replicate real-world scene distributions [56, 66, 75, 55, 20, 91] and are thus constrained by the complexity of modeling semantics, e.g. object affordances, our approach bypasses these challenges. This simplification enables highly scalable and efficient data generation.
@@
We generate the MegaSynth dataset, comprising 700K scenes. MegaSynth is over 50 times larger than the real dataset DL3DV and significantly scales up training data for LRMs. We utilize MegaSynth to train feed-forward LRMs [88, 93] jointly with DL3DV. Our experiments show a 1.2 to 1.8 dB PSNR gain across diverse test datasets and image resolutions. Moreover, the depth rendering quality is significantly improved, showing a better reconstruction geometry quality. These results highlight the synergy between synthesized and real data. Synthesized data excels in scale and provides rich metadata, such as geometry supervision, enabling models to develop a general geometric understanding beyond rendering supervision. Meanwhile, small-scale real data further sharpens the model. Interestingly, MegaSynth can also benefit other 3D tasks, where a monocular depth estimation model fine-tuned on MegaSynth demonstrates significant improvement.
@@ 4 Synthesizing the MegaSynth Dataset
Overview. We synthesize MegaSynth using a procedural generation method, as illustrated in Fig. 2. The process involves: i) generating a scene floor plan, including scene size and object instance box locations, ii) instantiating object geometries with random textures, and iii) randomizing the lighting. During the process, we eliminate high-level scene semantics. We only keep the low-level structural and geometric features of scenes.
@@ 4.1 Scene Floor Plan
Without loss of generality, we plan the scene as a cube box and populate it with objects represented by 3D bounding boxes. We randomize the 3D aspect ratio and size of scenes. We design multiple object box categories to simulate real-world scene geometry structures (visualized as boxes in Fig. 2 with different colors). For example, large object boxes tend to be placed near the ground while small object boxes have more flexible placement options. We parameterize the size, location, and number of each object type, specifying each parameter as a range. This allows us to introduce randomness to improve diversity. Further details of the object box categories and their attribute sampling ranges are provided in the Appendix.
@@ 4.3 Lighting
Luminous objects and light bulbs. We randomly turn objects and axis-aligned sticks as lights and place light bulbs in the scene, simulating real-world lighting and increasing diversity. The intensity of object light can be sampled as large values to simulate lighting in dark environments.
@@ 5 Learning 3D Reconstruction on MegaSynth
In this section, we discuss how we utilize our synthesized MegaSynth (Sec. 4) to train a feed-forward reconstruction model (i.e., the LRM-based model illustrated in Sec. 3). To reach the goal, we first construct the training data by carefully sampling the camera distribution and rendering the images (Sec. 5.1). We then train our model with a mixed-data training strategy (Sec. 5.2) with rendering loss and geometry loss (Sec. 5.3). The details of the training process can be found in the Appendix.
@@ 6 Experiments
Model Training Dataset PSNR SSIM LPIPS PSNR SSIM LPIPS AbsRel $\delta_{1\uparrow}$ PSNR SSIM LPIPS
@@
3DGS [37] N.A. (Per-scene Optimization) 5.2 min 24.27 0.817 0.166 20.67 0.672 0.293 0.320 0.715 16.46 0.458 0.405
Long-LRM [93] DL3DV 0.12 sec 24.18 0.812 0.173 23.41 0.790 0.210 0.272 0.763 19.68 0.569 0.312
Long-LRM (ours) DL3DV + MegaSynth 25.44 0.853 0.136 25.01 0.836 0.164 0.258 0.792 20.86 0.652 0.249
GS-LRM [88] DL3DV 0.11 sec 24.60 0.824 0.161 23.89 0.806 0.195 0.291 0.772 19.93 0.601 0.289
GS-LRM (ours) DL3DV + MegaSynth 25.75 0.859 0.130 25.46 0.846 0.154 0.258 0.800 21.19 0.672 0.235
@@
3DGS [37] N.A. (Per-scene Optimization) 6.4 min 23.26 0.778 0.206 21.75 0.690 0.294 0.319 0.709 16.06 0.436 0.421
Long-LRM [93] DL3DV 0.35 sec 23.71 0.779 0.236 22.51 0.767 0.267 0.291 0.753 18.61 0.465 0.421
Long-LRM (ours) DL3DV + MegaSynth 25.14 0.828 0.186 24.26 0.817 0.210 0.255 0.794 19.84 0.555 0.339
Table 1: Evaluation results against baseline methods. We report results at resolutions of 128 and 256. For resolution 256, we only report results of Long-LRM as transformer-based GS-LRM is too slow. Our models are pre-trained on MegaSynth and then tuned on DL3DV. We report NVS quality on all data and evaluate reconstruction by measuring geometry accuracy (rendered depth accuracy) on Hypersim.
In this section, we describe the experimental setting and present evaluation results. Due to the space limit, implementation and training details are in the Appendix.
@@
DL3DV-Test-Indoor
Data PSNR SSIM LPIPS
DL3DV 25.41 0.853 0.150
DL3DV + MegaSynth 26.75 0.890 0.116
DL3DV-Test-Outdoor
Data PSNR SSIM LPIPS
DL3DV 23.09 0.771 0.183
DL3DV + MegaSynth 23.89 0.803 0.157
Table 3: Performance gains on indoor and outdoor test data. Results are from 128-resolution GS-LRM. Test data split details are in Sec. 6.1.
@@ 6.1 Datasets
Hypersim [56] is a synthetic 3D indoor scene dataset with ultra photo-realistic renderings, aimed at testing the generalization capability to out-of-distribution indoor scenes. Hypersim is challenging due to its complicated geometry, extreme lighting conditions, and large camera baselines. Hypersim also provides high-quality depth ground truth. We use a test set composed of 302 scenes.
@@ 6.2 Evaluation and Baselines
We use $32$ views as input and use $32$ target novel view images for evaluation. The input and target views are non-overlap and are evenly sampled. We compare with three baselines:
@@
Additionally, we note that comparing with more advanced 3DGS methods is not the focus our work, as our target is scaling up training data for improving feed-forward methods. Our contributions can be directly ablated by comparing with LRMs trained without our data.
@@ 6.3 Results
Table 1 presents our results, demonstrating that training with both DL3DV and our MegaSynth dataset significantly improves model performance, with PSNR gains ranging from 1.2 to 1.8 dB. This improvement is consistent across model architectures (GS-LRM and Long-LRM), testing data (both in-domain DL3DV and out-of-domain datasets), image resolutions, and evaluation metrics, highlighting the effectiveness of our synthesized MegaSynth in enhancing the reconstruction quality of LRMs for wide-coverage scenes. Moreover, the rendering depth quality improves significantly as evaluated on Hypersim, showing the benefit of improving geometry quality by training with MegaSynth. Fig. 3 and Fig. 4 visually compare the reconstruction results for models with and without MegaSynth. We observe remarkable improvements in scenes with complicated scene structures, geoemtry, material and lighting, aligning with data generation designs (Sec. 4). Our approach also achieves substantially better results than the optimization-based 3DGS method while offering much faster inference speeds (e.g., over 2000 times speed-up from 5 minutes to 0.1 seconds).
@@ 6.4 Ablation Studies
In this section, we examine the impact of MegaSynth data quality, quantity, properties, and training paradigms for utilizing synthesized data. Without additional specification, the default experimental setup is the resolution-128 GS-LRM with pre-training + fine-tuning training protocol.
@@
Hypersim
@@
DL3DV Real-only 23.89 0.806 0.195
MegaSynth-only 21.50 0.719 0.272
Joint Training 25.33 0.844 0.157
Pre-training + Fine-tuning 25.46 0.846 0.154
Table 4: Ablation study on the training framework to leverage MegaSynth. Results are reported with GS-LRM.
Table 5: Visual comparison between training with only MegaSynth and only real data. We include two failure cases of MegaSynth-only with failures highlighted.
Table 6: Analysis of real data only and MegaSynth-only performance with different number of input views.
@@ 6.5 Analysis
We perform a more detailed analysis of MegaSynth, especially its effectiveness against other synthetic data and application to other 3D tasks.
@@
DL3DV Hypersim MipNeRF360 & TT
PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
DL3DV 18.31 0.555 0.391 18.43 0.602 0.373 15.59 0.550 0.332
DL3DV+Kubric [23] 18.28 0.552 0.395 18.46 0.600 0.375 15.49 0.546 0.340
DL3DV+Front3D [20] 18.40 0.558 0.389 18.48 0.603 0.370 15.63 0.551 0.329
DL3DV+MegaSynth 19.58 0.592 0.338 19.88 0.638 0.324 16.72 0.592 0.303
Table 7: Comparison with other synthetic datasets. We report results with 8 input views and GS-LRM under resolution 128.
AbsRel ($\downarrow$) $\delta_{1}(\uparrow)$
Depth Anything V2 0.213 0.761
Tuned on MegaSynth 0.158 0.799
Table 8: MegaSynth benefits monocular depth estimation.
DL3DV Ours Front3D
Geom. Difficulty ($\downarrow$) 1.65 1.35 3.00
Diversity ($\downarrow$) 1.40 1.60 3.00
Table 9: User study of data difficulty and diversity.
Advantages over other synthetic datasets. We experiment with using other synthetic datasets for training LRMs. As shown in Table 7, Kurbic [23] (data released in SRT [58]) and Front3D [20] fail to improve LRM performance, while MegaSynth benefits the model across all test datasets consistently. In detail, Kurbic contains 1 million scenes radnomly composed by realistic 3D assets; Front3D is composed of 6,000 indoor scenes designed by artists. The results imply that realistic 3D assets or scene composition is not the gaurantee for improving reconstruction quality. Instead, reconstruction model benefits from data with better non-semantic quality, e.g. geometry difficulty and scene diversity.
MegaSynth also helps other tasks. We fine-tune Depth Anything V2 ViT-B model on MegaSynth and evaluate on Hypersim. Results in Table 9 shows that MegaSynth helps improving monocular depth estimation, demonstrating the potential of MegaSynth to be used for other 3D tasks.
Comparison with real data. Tab. 9 presents a user study ranking geometry difficulty and scene diversity of datasets, showing our comparability with real data and advantage over the other synthetic data Front3D. Please see more analysis on measuring alignment with real data in Appendix.